OpenStudy (anonymous):

In math, how come 2(x+2) + 3(x+2) = (2+3)(x+2)

OpenStudy (shubhamsrg):

go backwards i.e. from RHS to LHS and see if you can understand..

OpenStudy (anonymous):

think of it as like terms the things in the parenthesis on the left are exactly alike.. so you're adding like terms...

OpenStudy (anonymous):

\[R.H.S=(2+3)(x+2)=2x+4+3x+6=2(x+2)+3(x+2)\]

OpenStudy (anonymous):

which is our R.H.S

OpenStudy (anonymous):

What is the math rule that allows this?

OpenStudy (anonymous):

@shubhamsrg @Aamal @dpaInc

OpenStudy (anonymous):

to obtain the desired result we take any side of our given equation and solve it we simply here multiply the two terms